Circular No. 6253

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995aj IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur (OCA), reports his discovery

of a supernova at α = 2^{h}28^{m}36^{s}.64, δ = +41^{o}56′01′′.9 (equinox 2000.0), which is 5′′.0 west and 4′′.9 south of the spiral galaxy’s center (the galaxy having mag B ∼ 18.5). Estimated magnitudes for SN 1995aj from OCA Schmidt telescope films taken by D. Albanese and Pollas: Sept. 29.06 UT, V ∼ 20.0; Oct. 25.00, 21.4. A nearby star (V ∼ 18.4) has position end figures 36^{s}.64, 21′′.3. Nothing is present at this location on the original Palomar Sky Survey prints or the 1989 (B_{j}) Palomar Survey film.

L. Wang, Department of Astronomy, University of Texas, communicates: “A low-resolution (1.2-nm FWHM) spectrogram of SN 1995aj was obtained on Oct. 26 UT with the McDonald Observatory’s 2.1-m telescope. Preliminary inspection of the spectrum shows this to be a type-Ia supernova ∼ 1 month past maximum.”

51 PEGASI N. Houk, University of Michigan, has determined a spectral type of

G2–3 V for 51 Peg from a 1-min IIa-O 10^{o} objective-prism plate taken on 1981 Aug. 6 with the Burrell Schmidt telescope at Kitt Peak. The high-quality spectrum (resolution ∼ 0.1 nm; dispersion 10.8 nm/mm, i.e., normal classification dispersion), together with the use of several standard spectra, revealed no duplicity or chemical peculiarities. The spectrum is definitely not as luminous as IV (quoted in the 1982 Bright Star Catalogue from Keenan and Pitts 1980, Ap.J. Suppl. 42, 561), or even ‘IV-V’. It is also earlier than the temperature type of G5 given in several older publications.”

Corrigendum. On IAUC 6251, line 24, for 0.037 read 0.0037

SUPERNOVA 1995ag IN UGC 11861 Magnitude and precise position obtained by E. Jung, Marl, Germany,

with a 0.20-m reflector (+ CCD) on Oct. 21.840 UT: R = 16.0, α = 21^{h}56^{m}19^{s}.5, δ = +73^{o}15′14′′.3 (equinox 2000.0).

COMET 29P/SCHWASSMANN-WACHMANN 1 Total CCD magnitudes: Oct. 20.80 UT, 13.4 (T. Kojima, Chiyoda,

Japan, 0.25-m reflector); 21.17, 13.2 (S. Garro, Hautes Alpes, France, 0.20m reflector; coma diameter 1′.8).

1995 October 26 Daniel W. E. Green